### v2

@@ 6 Discussion and Limitation
Our ValueCompass framework has revealed critical insights into human-AI value alignment across diverse contexts. The moderate alignment rates (highest F1 score of only 0.529) indicate substantial room for improvement, with notable variations across countries and scenarios. Humans frequently endorse values like ”National Security” that LLMs largely reject, while alignment exists on values such as ”Successful” and ”Capable.” Qualitative analysis further revealed that humans prioritize AI systems that remain subordinate to human control, maintain objectivity, avoid harm, and uphold principles like fairness.
@@
These findings highlight several important implications for AI development and governance. The contextual variations in alignment underscore the need for context-aware strategies rather than one-size-fits-all approaches. Many participants emphasized maintaining human oversight in AI-assisted decision-making, suggesting technical solutions should complement rather than replace human judgment. The identification of specific value misalignments suggests AI developers need explicit frameworks for prioritizing certain values in contexts where conflicts emerge. The ValueCompass framework offers a practical diagnostic tool to identify potential misalignments before deployment, potentially reducing ethical risks in production systems.
Limitations. Despite these contributions, several limitations must be acknowledged. Our human survey sample (112 participants across seven countries) may not fully capture global value diversity, and self-reported values may be subject to social desirability bias. Our LLM evaluation approach assumes models can accurately report their inherent values through prompted responses, potentially missing complex value encodings. Additionally, our study is limited in scenario coverage, focuses primarily on Western cultural contexts, captures values only at a static point in time, and relies on Schwartz’s theory which may not capture all AI-relevant value dimensions. Future work should address these limitations to develop more comprehensive evaluations of human-AI value alignment across diverse contexts.
@@ 7 Conclusion
In this work, we introduced ValueCompass, a comprehensive framework for identifying and evaluating human-AI alignment based on fundamental values derived from psychological theory and a systematic review. By applying ValueCompass across four real-world vignettes—collaborative writing, education, public sectors, and healthcare—we uncovered significant misalignments between human and language model (LM) values. Notably, humans frequently endorse values like ”National Security” which were largely rejected by LLMs. Our findings also demonstrated that value preferences vary across different contexts, underscoring the importance of developing context-aware alignment strategies for AI systems. This research provides crucial insights into the complexities of aligning AI systems with diverse human values, offering a foundational step toward creating AI technologies that responsibly and effectively reflect societal ethics and principles.
@@ A.1 Cultural and Social Values
We introduce the 56 universal values and their definitions outlined in the Schwartz’s Theory of Basic Values (Schwartz, 1994; 2012), which consists of 56 exemplary values covering ten motivational types. We show the complete list of value in Table 3.
@@ A.2 Prompt Variation Design
[…]

### v3

@@ 5 Discussion and Implications
Our ValueCompass framework has revealed critical insights into human-AI value alignment across diverse contexts. The moderate alignment rates (highest F1 score of only 0.529) indicate substantial room for improving value alignment, with notable variations across countries and scenarios. Humans frequently endorse values like “National Security” that LLMs largely reject, while alignment exists on values such as “Successful” and “Capable.” Qualitative analysis further revealed that humans prioritize AI systems that remain subordinate to human control, maintain objectivity, avoid harm, and uphold principles like fairness.
@@
Implications. These findings highlight several important implications for AI development and governance. The contextual variations in alignment underscore the need for context-aware strategies rather than one-size-fits-all approaches. Many participants emphasized maintaining human oversight in AI-assisted decision-making, suggesting technical solutions should complement rather than replace human judgment. The identification of specific value misalignments suggests AI developers need explicit frameworks for prioritizing certain values in contexts where conflicts emerge. The ValueCompass framework offers a practical diagnostic tool to identify potential misalignments before deployment, potentially reducing ethical risks in production systems.
6 Related Work
Evaluating LLM Values. Early studies focused on specific values such as  (Shen et al., 2022), interpretability (Shen et al., 2023), and safety (Zhang et al., 2020). Recent work has expanded to broader ethical frameworks (Kirk et al., 2024; Jiang et al., 2024; Sorensen et al., 2024), often using fixed datasets like the World Value Survey (Haerpfer et al., 2020). However, these approaches lack generalizability. Others use limited value sets from Moral Foundations Theory (Park et al., 2024), which miss dimensions like honesty and creativity. In contrast, our work applies Schwartz’s Theory of Basic Values (Schwartz, 1994, 2012) for a broader, cross-cultural evaluation across contexts.
Human–AI Value Alignment. Most prior work treats alignment as part of AI safety, focusing on model-side alignment (Dillion et al., 2023). Recent studies consider human–AI bidirectional-alignment Shen et al. (2024) and use prompt-based evaluations (Norhashim and Hahn, 2024), but lack a generalizable framework. We address this gap by systematically evaluating human–LLM alignment across diverse values and scenarios.
@@ 7 Conclusion
We introduced ValueCompass, a framework for evaluating human–AI alignment using fundamental values from psychological theory. Applied to four real-world contexts—collaborative writing, education, public sectors, and healthcare—it revealed significant misalignments, such as LLMs rejecting values like National Security that humans frequently endorse. Our results highlight the need for context-aware alignment strategies and offer a foundation for developing AI systems that better reflect human values and societal principles.
Limitations
Despite these contributions, several limitations must be acknowledged. Our human survey sample (112 participants across seven countries) may not fully capture global value diversity, and self-reported values may be subject to social desirability bias. Our LLM evaluation approach assumes models can accurately report their inherent values through prompted responses, potentially missing complex value encodings. Additionally, our study is limited in scenario coverage, focuses primarily on Western cultural contexts, captures values only at a static point in time, and relies on Schwartz’s theory which may not capture all AI-relevant value dimensions. Future work should address these limitations to develop more comprehensive evaluations of value alignment across diverse contexts.
Acknowledgement
We sincerely thank Michael Terry for his valuable insights and contributions, and Meredith Ringel Morris for her thoughtful review and encouraging feedback. We greatly appreciate Matías Duarte for his support and constructive comments, and Savvas Petridis for his review and help. Finally, we thank all participants of the human survey studies for their contributions. This project was partly supported by the National Science Foundation under Grant No. 2119589 and by the Institute of Museum and Library Services RE-252329-OLS-22.
@@ Appendix A Appendix
Figure 3: Value Form is a context-aware instrument to measure the value alignment between humans and LLMs. It includes a task introduction, a vignette, and 56 value statements, grounded in Schwartz Theory of Basic Values. As shown in Figure 1, humans and LLMs rate each value on a scale from “-2: Strongly Disagree” to “2: Strongly Agree”, plus “Irrelevant." The form aims to assess human-AI value alignment contextualized in various scenarios.
@@ A.1 Cultural and Social Values
We introduce the 56 universal values and their definitions outlined in the Schwartz’s Theory of Basic Values (Schwartz, 1994, 2012), which consists of 56 exemplary values covering ten motivational types. We show the complete list of value in Table 3.
@@ A.2 Prompt Variation Design
Figure 4: Four vignettes, designed to contextualize the value statements in the ValueCompass framework, are organized by increasing risk and reflect real-world tasks: collaborative writing, education, the public sector, and healthcare. Images are included in the vignettes to aid respondents in understanding the context.